# CATArena: Evaluation of LLM Agents Through Iterative Tournament Competitions

### v1

@@ Appendix I Results of Multi-lingual Track
Average across games
Total input tokens 1277124.75 53501.75 385533.5 599261.75 741507.75 1478472.25 1144616.25 - 597675.0 876130.5
Complete tokens 93389.25 11325.25 70819.5 76283.5 57610.5 42925.75 43872.0 - 39733.25 39306.5
Total output tokens 13538.5 1423.25 16999.0 28075.5 8760.0 14102.5 14683.0 - 8596.75 8802.0
Total time (s) 421.4060 35.8547 363.4863 553.4572 511.2015 7458.3587 480.9153 - 266.3653 441.8328
Tools used 49.5 15.5 21.5 49.5 54.0 209.75 37.75 20.75 19.0 33.0
Valid lines of code 575.5 39.5 400.25 239.25 318.0 283.25 412.25 321.75 267.5 403.0
Avg thinking time (s) 0.4141 0.0012 0.0247 0.0134 0.2137 0.0477 0.0258 0.0273 0.4330 0.0059
We list the agents’ cost and code statistics in table 15 for first round development of standard games,  16 for second round development of standard games, 17 for first round development of vairant games and  18 for second round development of variant games.
We can see that game development token costs show minimal variation, while differences is significant due to model changes. Claude (both minimal and code-based agents) consumes significantly more input tokens than competitors, exceeding the average by over 2 times, while Gemini generates notably more output tokens compared to other models. GPT-5 offers the best trade-off between cost. Among all agents, second-round game development require more input tokens, while output token growth remains marginal. In addition, commercial agents consistently use fewer tokens than their minimal-agent counterparts.
In terms of code complexity, agents driven by Claude-4 model consistently surpasses other agents in both the number of effective lines of code developed and the time spent considering development strategies. We observe that its development strategies are more sophisticated. Additionally, the complexity of its code increases with each iteration, which indirectly demonstrates the model’s exceptional learning capabilities.
Appendix K Full Prompts
The agent is instructed to develop a competitive game AI based on the provided game environment. The AI must be deployed as an HTTP service with a single-port startup script, follow the official development instructions, and be named with the model prefix. And the agent is encouraged to iteratively improve its strategy based on the tournament report and the previous codes. Full prompt of details are in Table 19 for main leaderboard,  LABEL:tab:ml_game_ai_prompt for ML track and  21 for multilingual track.
@@
The final AI should be provided as an HTTP service. You can refer to the guides in game env path / README.md and game env path / develop instruction.md for development instructions. The content in game env path / develop instruction.md is very important, please read it carefully!
@@
You can play game of game env pathin at game server. You can play the games with your own AI or any other AI to improve your strategy. You can use bash tools to improve yourself.
Final Remind
@@
Tournament report of last round is in last round log dir. The historical records are quite large, please use tools start_interactive_shell and run_interactive_shell to analyze the data efficiently. The code corresponding to the log is stored in last round code dir. Please learn from it and improve your strategy.
@@ 4. Training Endpoint: /train HTTP endpoint to trigger training
Table 21: Multi-language Game AI Development Prompt
@@
The final AI should be provided as an HTTP service. You can refer to the guides in game env path / README.md and game env path / develop instruction.md for development instructions. The content in game env path / develop instruction.md is very important, please read it carefully!
@@
You can play game of game env pathin at game server. You can play the games with your own AI or any other AI to improve your strategy. You can use bash tools to improve yourself.
Final Remind

### v2

@@ Appendix I Case Study of Agent Learning
OIBench
Total input tokens 1604531 40210 8364987 620122 665089 1081920 1027936 46854 - -
Complete tokens 1610748 42112 8493081 653688 704924 1095605 1033256 78278 - -
Total output tokens 6217 1902 128094 33566 39835 13685 5320 31424 - -
Total time (s) 214.430 60.390 3604.330 3708.590 1235.390 411.730 259.790 810.060 - -
Tools used 20 8 135 34 33 39 34 8 - -
Valid lines of code 256 43 391 396 262 222 241 0 600 357
Avg thinking time (s) - - - - - - - - - -
Appendix O Full Prompts
The agent is instructed to develop a competitive game AI based on the provided game environment. The AI must be deployed as an HTTP service with a single-port startup script, follow the official development instructions, and be named with the model prefix. The agent is encouraged to iteratively improve its strategy based on the tournament report and the previous solutions. Full prompt details are in Tables 18 for the main leaderboard,  LABEL:tab:ml_game_ai_prompt for ML track and  20 for multilingual track.
@@
You can play the game at game server using game env path. You can play the games with your own AI or any other AI to improve your strategy. You can use bash tools to improve yourself.
Final Reminder
@@
The tournament report from the last round is in last round log dir. The historical records are quite large, please use tools start_interactive_shell and run_interactive_shell to analyze the data efficiently. The code corresponding to the log is stored in last round code dir. Please learn from it and improve your strategy.
@@ 4. Training Endpoint: /train HTTP endpoint to trigger training
Table 20: Multilingual Game AI Development Prompt
@@
You can play the game at game server using game env path. You can play the games with your own AI or any other AI to improve your strategy. You can use bash tools to improve yourself.
Final Reminder